# Holographic Boundary Conformal Field Theory within Horndeski Gravity

### v1

@@ III Entanglement entropy and Hatman-Maldacena surface
III.1 Hartman-Maldacena entanglement entropy
@@
III.2 Field theory calculation
@@
IV Page curve behavior
@@
In Figs. 9 and 10 we depict the evolution of the Page angle density $\rho_{Page}$ as a function of $r_{L}/r_{h}$ and $r_{R}/r_{h}$, where $r_{L}$ and $r_{R}$ are the positions of the left and right boundaries, for various values of $\gamma$ and $\alpha$. These figures demonstrate that the Horndeski parameters can induce significant changes in the system, particularly when the radial coordinate $r$ is near 00 or the horizon radius $r_{h}$ Geng:2022dua . Such changes align with the understanding that Horndeski gravity causes a shift in the islands connected by the Hartman-Maldacena surface Hartman:2013qma . Our findings reveal that the residual information, denoted as $S^{L}_{bdry}$ and $S^{R}_{bdry}$, significantly affects the location of the bipartition, except when it is near the event horizon or a defect. Moreover, these figures illustrate the dependence of the Page curve behavior on the Horndeski parameters $\alpha$ and $\gamma$, as well as on the geometric factors $r_{L}$ and $r_{R}$. Finally, they provide visual evidence of the significant role that Horndeski gravity plays in modifying the dynamics of entanglement entropy compared to standard general relativity.
@@
In Figs. 11 and 12 we present the evolution of the density of the Page angle $\rho_{Page}$ about $r_{L}$ and $r_{R}$, for various values of $\gamma$ and $\alpha$. The horizontal axis represents $r/r_{h}$, ranging from 0 to 1, the vertical axis shows the Page angle $\rho_{\star}$, and the color gradient indicates the logarithm of the Page time. The white dotted curves depict the contour where $\rho_{Page}=0$, which represents the transition point between island dominance ($\rho_{Page}>0$) and HM surface dominance ($\rho_{Page}<0$) at $t=0$. As $\gamma$ changes from $-1$ to $-0.1$, we observe significant changes in the density patterns, particularly near the corners of the plot where $r_{L}/r_{h}$ or $r_{R}/r_{h}$ approach 00 or $1$. These changes demonstrate the strong influence of the Horndeski parameter $\gamma$ on the Page angle, especially in regions close to the defect or the horizon. Regions where $\rho_{\star}>0$ indicate island dominance at $t=0$, while $\rho_{\star}<0$ shows Hartman-Maldacena surface dominance. As $\gamma$ approaches zero, we observe more pronounced variations in the Page angle, particularly for smaller values of $r/r_{h}$. Additionally, as $\alpha$ changes from $-11/3$ to $-1$, we observe significant changes in the density patterns, particularly in regions where $r_{L}/r_{h}$ or $r_{R}/r_{h}$ are close to 00 or $1$.

### v2

@@ 3.2 Derivation of the functional entanglement entropy for Horndeski gravity
3.3 Hartman-Maldacena entanglement entropy
@@
3.4 Field theory calculation
@@
4 Page curve behavior
@@
In Figs. 12 and 13 we depict the evolution of the Page angle density $\rho_{Page}$ as a function of $r_{L}/r_{h}$ and $r_{R}/r_{h}$, where $r_{L}$ and $r_{R}$ are the positions of the left and right boundaries, for various values of $\gamma$ and $\alpha$. These figures demonstrate that the Horndeski parameters can induce significant changes in the system, particularly when the radial coordinate $r$ is near 00 or the horizon radius $r_{h}$ Geng:2022dua . Such changes align with the understanding that Horndeski gravity causes a shift in the islands connected by the Hartman-Maldacena surface Hartman:2013qma . Our findings reveal that the residual information, denoted as $S^{L}_{bdry}$ and $S^{R}_{bdry}$, significantly affects the location of the bipartition, except when it is near the event horizon or a defect. Moreover, these figures illustrate the dependence of the Page curve behavior on the Horndeski parameters $\alpha$ and $\gamma$, as well as on the geometric factors $r_{L}$ and $r_{R}$. Finally, they provide visual evidence of the significant role that Horndeski gravity plays in modifying the dynamics of entanglement entropy compared to standard general relativity.
@@
In Figs. 14 and 15 we present the evolution of the density of the Page angle $\rho_{Page}$ about $r_{L}$ and $r_{R}$, for various values of $\gamma$ and $\alpha$. The horizontal axis represents $r/r_{h}$, ranging from 0 to 1, the vertical axis shows the Page angle $\rho_{\star}$, and the color gradient indicates the logarithm of the Page time. The white dotted curves depict the contour where $\rho_{Page}=0$, which represents the transition point between island dominance ($\rho_{Page}>0$) and HM surface dominance ($\rho_{Page}<0$) at $t=0$. As $\gamma$ changes from $-1$ to $-0.1$, we observe significant changes in the density patterns, particularly near the corners of the plot where $r_{L}/r_{h}$ or $r_{R}/r_{h}$ approach 00 or $1$. These changes demonstrate the strong influence of the Horndeski parameter $\gamma$ on the Page angle, especially in regions close to the defect or the horizon. Regions where $\rho_{\star}>0$ indicate island dominance at $t=0$, while $\rho_{\star}<0$ shows Hartman-Maldacena surface dominance. As $\gamma$ approaches zero, we observe more pronounced variations in the Page angle, particularly for smaller values of $r/r_{h}$. Additionally, as $\alpha$ changes from $-11/3$ to $-1$, we observe significant changes in the density patterns, particularly in regions where $r_{L}/r_{h}$ or $r_{R}/r_{h}$ are close to 00 or $1$.